Please subscribe to the ACO-status mailing list for future ACO operational advisories and alerts.

The ACO website will be down on June 28, 2018 due to scheduled network maintenance.

New instruments were installed during the ACO-6 cruise on 6/20/2018 (hydrophone, secondary node). BSP1 and BSP2 disconnected from ACO until further notice. Data affected: CTD2, CTD3, Pressure and Fluorometer.

ALOHA Cabled Observatory
Hydrophone Experiment Module Engineering Data
Temperature
Temperature 1 :
Temperature 2 :
Temperature 3 :
Temperature 4 :
12.69929
12.883644
15.335172
8.681387
°C
°C
°C
°C
Date: July 17, 2018    13:25:36 UTC
Pressure, Current, Voltage
Pressure:
Center +48 V dc:
+48 V dc :
Ground Fault Offset:
Voltage Reference:
3.20048
24.024
48.144
-0.096
2.487
PSI
V
V
V
V
Date: July 17, 2018    13:25:36 UTC